# A variational formulation of stochastic thermodynamics. Part I: Finite-dimensional systems

### v1

@@ V Conclusion
In summary, the results presented here show that the variational formulation of stochastic thermodynamics provides a unified and systematic geometric framework for constructing models that are thermodynamically consistent by design. Crucially, this approach remains valid far from equilibrium, applies to both closed and open systems, and naturally accommodates nonlinear system-bath interactions beyond weak coupling. It systematically integrates microscopic fluctuations with macroscopic thermodynamic structure, offering a flexible and modular formalism grounded in the second law. These features make it broadly applicable, from isolated mechanical systems to open, interacting, and actively driven environments.
Remarkably, fluctuation-dissipation relations have emerged as fundamental constraints for thermodynamic consistency beyond detailed balance. In particular, they have been shown to ensure compliance with the second law of thermodynamics and to provide a well-defined interpretation of entropy production. Within this context, the variational approach provides a systematic framework for deriving such relations from first principles.
Ongoing work extends this framework to infinite-dimensional systems (fields), where the underlying geometric structure of the variational principle enables thermodynamically consistent modeling of complex fluids, active matter, and other continuum systems. Promising research directions include the development of variational integrators for structure-preserving numerical schemes [64, 65, 66], the application of Lagrangian reduction by symmetry to capture emergent phenomena [67, 68], and the incorporation of non-Markovian dynamics within this formalism.
@@ Acknowledgements.
H. V. del P. acknowledges the Provost Graduate Award scholarship from Nanyang Technological University, Singapore, and thanks J. M. R. Parrondo for valuable discussions on information thermodynamics. F. G.-B. is partially supported by a startup grant from Nanyang Technological University and by the Ministry of Education, Singapore, under Academic Research Fund (AcRF) Tier 1 Grant RG99/24. H.Y. is partially supported by JSPS Grant-in-Aid for Scientific Research (22K03443), JST CREST (JPMJCR24Q5), and Waseda University (SR 2025C-095). H. V. del P. derived the results and prepared the original draft. F. G.-B. and L. Y. C. supervised the work. F. G.-B., H. Y., and L. Y. C. contributed to the review and editing of the manuscript. All authors participated in the discussions.
@@ Appendix A Two-step average proof
where the last equality is given by definition (3), and the proof is completed.
Appendix B Mechanical isolated system derivations
@@
In this section the detailed derivations of Sec. III.3 are provided. First, we derive the EoMs (LABEL:eq:EoM_mech_closed). The variational principle in this case is of the form:
@@
The total EPR then yields:
$\dot{S}_{tot}$ $=\dot{\mathcal{S}}+\langle\!\langle\dot{s}\rangle\!\rangle=\int d\Omega\,\frac{{j}_{d}^{2}}{\mathcal{P}g^{2}}$ (113)
$-\int d\Omega\,{j}_{d}\left\{\frac{p}{mT}+\frac{1}{g^{2}}\left(-\gamma\frac{p}{m}-g\partial_{p}g+\frac{p}{m}g\partial_{s}(g/T)\right)\right\},$
and the FDR (49) is found by setting to zero the expression enclosed by the brackets $\{$...\}.
Appendix C Interconnected systems derivations
In this section the detailed derivations of Sec. III.4 are provided. Firstly, taking variations of the Lagrangian and taking into account the fixed endpoints condition gives
@@ Appendix C Interconnected systems derivations
$\delta W^{i}:\dot{N}_{i}=(\mathcal{J}_{j\rightarrow i}+\sigma_{j\rightarrow i}\circ\zeta),$ (115d)
@@
$=\int\frac{d\Omega}{(1-C^{2})}\left\{\frac{j^{2}_{N_{1}}}{P\sigma^{2}}+\frac{j^{2}_{th}}{P\kappa^{2}}+2C\frac{j_{N_{1}}j_{th}}{P\sigma\kappa}+j_{N_{1}}\left(\frac{\tilde{\mathcal{J}}_{12}}{\sigma^{2}}-\frac{C}{\kappa\sigma}\tilde{J}_{12}\right)-j_{th}\left(\frac{\tilde{J}_{12}}{\kappa^{2}}-\frac{C}{\kappa\sigma}\tilde{\mathcal{J}}_{12}\right)\right\},$ (119)
where we used Eqs. (67LABEL:eq:j_th_inter). Adding both contributions and recalling the matrices $\mathbb{L},\mathbb{L}^{-1}$ as defined in (73), the average total EPR gives the form, with the shorthand notation $\bm{\mathfrak{J}}=(j_{N_{2}},j_{th})$:
$\dot{S}_{tot}$ $=\int\frac{d\Omega}{\mathcal{P}}\bm{\mathfrak{J}}^{T}\mathbb{L}^{-1}\bm{\mathfrak{J}}$
$-\int d\Omega\,\bm{\mathfrak{J}}\cdot\left\{\mathbb{L}^{-1}\begin{bmatrix}\tilde{\mathcal{J}}_{12}\\ \tilde{J}_{12}\end{bmatrix}-\begin{bmatrix}\vskip 5.69054pt\displaystyle\frac{\mu^{1}}{T^{1}}-\frac{\mu^{2}}{T^{2}}\\ \displaystyle\frac{T^{1}-T^{2}}{T^{1}}+\frac{T^{2}-T^{1}}{T^{2}}\end{bmatrix}\right\}$. (120)
The FDR (72) is then found by setting to zero the expression enclosed by the brackets $\{$...\}. The matrix $\mathbb{L}$ has eigenvalues
which satisfy $\lambda_{\pm}>0$ for $|C|<1$, ensuring that $\mathbb{L}$ is positive definite. In this case, its inverse $\mathbb{L}^{-1}$ is also positive definite, since its eigenvalues are $1/\lambda_{\pm}>0$, which guarantees that $\dot{S}_{tot}\geq 0$. Recall that the extreme case $C=\pm 1$ render the model ill-defined due to redundancies.
Appendix D Mechanical open system derivations
In this section the detailed derivations of Sec. III.5 are provided. First, we derive the EoMs (LABEL:eq:EoM_open). The variational principle in this case is of the form:

### v2

@@ V Conclusion
The variational formulation of ST provides a unified and systematic geometric framework for constructing models that are thermodynamically consistent by design. Crucially, this approach remains valid far from equilibrium, applies to both closed and open systems, and naturally accommodates nonlinear bath interactions through coupling of configurational and thermal variables. It systematically integrates microscopic fluctuations with macroscopic thermodynamic structure, offering a flexible and modular formalism grounded in the second law. These features make it broadly applicable, from isolated mechanical systems to open, interacting, and actively driven environments.
Remarkably, FDRs have emerged as fundamental constraints for thermodynamic consistency to ensure compliance with the second law of thermodynamics, providing a well-defined interpretation of entropy production and ensuring LDB. Within this context, the variational approach offers a systematic framework to derive such relations from first principles, under the assumptions of time-scale separation, large bath size and micro-reversibility. Although we have restricted our results to Gaussian noise, we expect the nonholonomic constraints to be independent of the particular noise correlations. Exploring this direction—by extending the formalism to incorporate non-Markovian or non-Gaussian dynamics, for instance through generalized FP equations [24]—remains an interesting avenue for future research. Finally, as a variational formulation rooted in Hamilton’s principle, the current framework naturally applies to Langevin-type dynamics, but not to more general Markov chains.
Ongoing work extends this framework to infinite-dimensional systems (fields), where the underlying geometric structure of the variational principle enables thermodynamically consistent modeling of complex fluids, active matter, and other continuum systems. Promising research directions include the development of variational integrators for structure-preserving numerical schemes [37, 38, 39], the application of Lagrangian reduction by symmetry to capture emergent phenomena [40, 41], a systematic treatment of conservation laws and Lie group symmetries, and, as previously mentioned, the role of configurational-thermal variables coupling and non-Markovian dynamics.
@@ Acknowledgements.
H. V. del P. acknowledges the Provost Graduate Award scholarship from Nanyang Technological University, Singapore, and thanks J. M. R. Parrondo for valuable discussions on information thermodynamics. F. G.-B. is partially supported by a startup grant from Nanyang Technological University and by the Ministry of Education, Singapore, under Academic Research Fund (AcRF) Tier 1 Grant RG99/24. H.Y. is partially supported by JSPS Grant-in-Aid for Scientific Research (22K03443), JST CREST (JPMJCR24Q5), and Waseda University (SR 2025C-095). The authors thank the Referees for their careful review and constructive suggestions, which have improved the manuscript. H. V. del P. derived the results and prepared the original draft. F. G.-B. and L. Y. C. supervised the work. F. G.-B., H. Y., and L. Y. C. contributed to the review and editing of the manuscript. All authors participated in the discussions.
@@ Appendix A Two-step average proof
where the last equality is given by definition (II.1), and the proof is completed.
Appendix B Positivity of entropy production functional
Recall from the original work by Kullback and Liebler [54] that the KL-divergence is the functional of probability measures that satisfies [18]:
  1. 1.
    Non-negativity: $D_{KL}[P\|P^{\dagger}]\geq 0$, consistent with the second law of thermodynamics when interpreted as entropy production, and $D_{KL}[P\|P^{\dagger}]=0\iff P=P^{\dagger}$, consistent with thermodynamic equilibrium.
  2. 2.
    Strong additivity: ensuring consistency under composition of processes, and yielding extensivity of entropy in the case of independent subsystems.
  3. 3.
    Invariance: The divergence is invariant under non-singular phase space coordinate transformations (diffeomorphisms), ensuring EP is an objective physical observable.
Among information divergences, the KL-divergence is the unique functional (up to a positive multiplicative constant) that simultaneously satisfies non-negativity, strong additivity (the chain rule), and invariance, once pathological solutions are excluded by mild regularity assumptions (e.g. continuity or measurability). For axiomatic foundations and related characterizations of the KL-divergence, see the Shore-Johnson axioms [83], and modern treatments [58, 18, 85].
Let us now recover the expression for the total EP:
$\langle\Delta s_{\rm tot}\rangle$ $=\langle\Delta s_{sys}\rangle+\langle\Delta\Sigma\rangle=\langle\Delta{\mathfrak{s}}_{\rm tot}\rangle-\langle\Delta{\mathfrak{s}}_{m}\rangle+\langle\Delta\Sigma\rangle$ (118)
$=\langle\Delta{\mathfrak{s}}_{\rm tot}\rangle-\int_{0,\Omega}^{t}d\tau\,d\mathbf{u}\>\bm{\mathfrak{j}}\cdot\bm{\phi},$
where we have applied $\Delta s_{sys}=\Delta{\mathfrak{s}}_{\rm tot}-\Delta{\mathfrak{s}}_{m}$ and exchanged the integration order assuming the integrand is absolutely integrable (Fubini’s theorem) as $\int_{\mathbf{u}_{0}}^{\mathbf{u}_{t}}\mathcal{D}\mathbf{u}P[\mathbf{u}]\int_{0}^{t}d\tau\,(\dot{\Sigma}-\dot{\mathfrak{s}}_{m})=\int_{0}^{t}d\tau\langle\!\langle\dot{\Sigma}-\dot{\mathfrak{s}}_{m}\rangle\!\rangle$. Here, $\bm{\mathfrak{j}}$ denotes the dissipative contributions to the probability currents and $\bm{\phi}(\mathbf{u},\tau)$ is a measure of LDB violation $(\text{LDB}\iff\bm{\phi}\equiv 0\iff\dot{\Sigma}=\dot{\mathfrak{s}}_{m})$ independent of $\mathcal{P}$. This form $\langle\!\langle\dot{\Sigma}-\dot{\mathfrak{s}}_{m}\rangle\!\rangle=\int_{\Omega}d\mathbf{u}\>\bm{\mathfrak{j}}\cdot\bm{\phi}$ is obtained by direct inspection on a case-by-case basis; see Eqs. (53), (III.5), and (98). Note here we assumed a continuous phase space $\Omega$, but the argument is completely analogous for discrete phase spaces, as well as for any system dimensionality.
Since $\langle\Delta\mathfrak{s}_{\rm tot}\rangle:=D_{KL}[P[\mathbf{u}]||P[\hat{\mathbf{u}}$]] is given as a KL-divergence between forward and backward trajectories $\mathbf{u}:[0,t]\subset\mathbb{R}\rightarrow\Omega$, it satisfies properties (13)(1-3). The second term satisfies strong additivity (2). This follows by noting that it can be written as $\langle\Delta\Sigma-\Delta{\mathfrak{s}}_{m}\rangle$, for which strong additivity can be established independently for each contribution.
For $\langle\Delta\mathfrak{s}_{m}\rangle=\int_{\mathbf{u}_{0}}^{\mathbf{u}_{t}}\mathcal{D}\mathbf{u}P[\mathbf{u}]\text{ln}\frac{P[\mathbf{u}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}|\mathbf{u}_{t}]}$, its strong additivity follows from factorization of conditional path measures:
$\langle\Delta\mathfrak{s}_{m}\rangle$ $=\int\mathcal{D}\mathbf{u}\,P[\mathbf{u}]\ln\frac{P[\mathbf{u}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}|\mathbf{u}_{t}]}=\int\mathcal{D}\mathbf{u}\,P[\mathbf{u}]\left[\ln\frac{P[\mathbf{u}_{A}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}_{A}|\mathbf{u}_{t}]}+\ln\frac{P[\mathbf{u}_{B}|\mathbf{u}_{A},\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}_{B}|\hat{\mathbf{u}}_{A},\mathbf{u}_{t}]}\right]$ (119)
$=\int\mathcal{D}\mathbf{u}_{A}\,P[\mathbf{u}_{A}]\ln\frac{P[\mathbf{u}_{A}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}_{A}|\mathbf{u}_{t}]}+\Big\langle\int\mathcal{D}\mathbf{u}_{B}\,P[\mathbf{u}_{B}|\mathbf{u}_{A}]\ln\frac{P[\mathbf{u}_{B}|\mathbf{u}_{A},\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}_{B}|\hat{\mathbf{u}}_{A},\mathbf{u}_{t}]}\Big\rangle_{\mathbf{u}_{A}},$
where Bayes theorem $P[\mathbf{u}_{A},\mathbf{u}_{B}|\mathbf{u}_{0}]=P[\mathbf{u}_{A}|\mathbf{u}_{0}]P[\mathbf{u}_{B}|\mathbf{u}_{A},\mathbf{u}_{0}]$ and normalization of conditional probabilities have been applied [50, 54].
Notably, the log ratio of path weights immediately implies invariance. Defining the Jacobian $\mathcal{M}(\mathbf{z})=|\det(\partial\mathbf{u}/\partial\mathbf{z})|$ [54]:
$\langle\Delta\mathfrak{s}_{m}\rangle$ $=\int\mathcal{D}\mathbf{u}\,P[\mathbf{u}]\ln\frac{P[\mathbf{u}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}|\mathbf{u}_{t}]}$ (120)
$=\int\mathcal{D}\mathbf{z}\,\mathcal{M}(\mathbf{z})\mathcal{M}^{-1}(\mathbf{z})P[\mathbf{z}]\ln\frac{P[\mathbf{z}|\mathbf{z}_{0}]\mathcal{M}^{-1}(\mathbf{z})}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]\mathcal{M}^{-1}(\mathbf{z})}$
$=\int\mathcal{D}\mathbf{z}\,P[\mathbf{z}]\ln\frac{P[\mathbf{z}|\mathbf{z}_{0}]}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]}=\langle\Delta\mathfrak{s}_{m}\rangle,$
which is also reflected by the scalar product in Eq. (118).
The same scalar product structure ensures the invariance of $\langle\Delta\Sigma\rangle$, arising due to its construction as an invariant in the nonholonomic constraints (39). Importantly, both $\langle\Delta\Sigma\rangle$ and $\langle\Delta\mathfrak{s}_{m}\rangle$ admit the same generic representation in terms of probability currents, i.e. $\int_{0,\Omega}^{t}d\tau\,d\mathbf{u}\>\bm{\mathfrak{j}}\cdot\bm{\psi}$, where $\bm{\psi}(\Omega)$ is a state-dependent function independent of the probability measure $\mathcal{P}$. This shared structure suggests that strong additivity extends to $\langle\Delta\Sigma\rangle$ as well, a fact that can be verified explicitly. In this representation, strong additivity leads to a splitting into marginal and conditional terms in the probability measure $\mathcal{P}$, as opposed to the path weight $P$, since the time integration has been exchanged.
For FP dynamics of the form $\partial_{t}\mathcal{P}=-\partial_{\nu}\mathfrak{j}^{\nu}$ with
$\mathfrak{j}^{\nu}=\mathfrak{j}^{\nu}_{\rm drift}-\mathfrak{j}^{\nu}_{\rm diff}=F^{\nu}_{\rm drift}\mathcal{P}-D^{\nu\mu}\mathcal{P}\partial_{\mu}\ln\mathcal{P},$ (121)
this property is already explicit at the level of the diffusive contribution:
$\int dxdy\,\mathfrak{j}^{\nu}_{\rm diff}(x,y)\psi_{\nu}(x,y)$ $=\int dxdy\,D^{\nu\mu}\mathcal{P}(x,y)\partial_{\mu}\ln\mathcal{P}(x,y)\psi_{\nu}(x,y)$ (122)
$=\int dxdy\,D^{\nu\mu}\psi_{\nu}(x,y)\mathcal{P}(x)\mathcal{P}(y|x)\partial_{\mu}\ln\mathcal{P}(x)+\int dxdy\,D^{\nu\mu}\psi_{\nu}(x,y)\mathcal{P}(x)\mathcal{P}(y|x)\partial_{\mu}\ln\mathcal{P}(y|x)$
$=\int dx\,D^{\nu\mu}\mathcal{P}(x)\partial_{\mu}\ln\mathcal{P}(x)\langle\psi_{\nu}(x,y)\rangle_{y|x}+\Big\langle\int dy\,D^{\nu\mu}\mathcal{P}(y|x)\partial_{\mu}\ln\mathcal{P}(y|x)\psi_{\nu}(x,y)\Big\rangle_{x}$
$=\int dx\,\mathfrak{j}^{\nu}_{\rm diff}(x)\langle\psi_{\nu}(x,y)\rangle_{y|x}+\Big\langle\int dy\,\mathfrak{j}^{\nu}_{\rm diff}(y|x)\psi_{\nu}(x,y)\Big\rangle_{x}.$
The drift contribution can be absorbed into either the marginal or the conditional part, as
$\int dxdy\,\mathfrak{j}^{\nu}_{\rm drift}(x,y)\psi_{\nu}(x,y)$ $=\int dxdy\,F^{\nu}_{\rm drift}(x,y)\mathcal{P}(x)\mathcal{P}(y|x)\psi_{\nu}(x,y)$ (123)
$=\int dx\,\mathcal{P}(x)\langle F^{\nu}_{\rm drift}(x,y)\psi_{\nu}(x,y)\rangle_{y|x}=\Big\langle\int dy\,\mathfrak{j}^{\nu}_{\rm drift}(y|x)\psi_{\nu}(x,y)\Big\rangle_{x}.$
Consequently, the drift term plays no role in testing strong additivity; the latter is entirely controlled by the diffusive current.
An analogous decomposition holds for Master equations, where the joint probability evolves as
$d_{t}\mathcal{P}(x,y)=\sum_{x^{\prime},y^{\prime}}\mathfrak{j}^{y,y^{\prime}}_{x,x^{\prime}}=\sum_{x^{\prime},y^{\prime}}[w^{y,y^{\prime}}_{x,x^{\prime}}\mathcal{P}(x^{\prime},y^{\prime})-w^{y^{\prime},y}_{x^{\prime},x}\mathcal{P}(x,y)],$ (124)
and $w^{y,y^{\prime}}_{x,x^{\prime}}$ denotes the transition rate $(x^{\prime},y^{\prime})\to(x,y)$. Using Bayes rule and the identity $\sum_{x^{\prime},y^{\prime}}\mathfrak{j}^{y,y^{\prime}}_{x,x^{\prime}}=\sum_{x^{\prime}}\mathfrak{j}^{y}_{x,x^{\prime}}+\sum_{y^{\prime}}\mathfrak{j}^{y,y^{\prime}}_{x}$ [48], one finds:
$\sum_{x,y}\sum_{x^{\prime},y^{\prime}}$ $\mathfrak{j}^{y,y^{\prime}}_{x,x^{\prime}}\psi^{y,y^{\prime}}_{x,x^{\prime}}=\sum_{x,y}\sum_{x^{\prime}}\mathfrak{j}^{y}_{x,x^{\prime}}\psi^{y}_{x,x^{\prime}}+\sum_{x,y}\sum_{y^{\prime}}\mathfrak{j}^{y,y^{\prime}}_{x}\psi^{y,y^{\prime}}_{x}$ (125)
$=\sum_{x,y}\sum_{x^{\prime}}[w^{y}_{x,x^{\prime}}\mathcal{P}(x^{\prime},y)-w^{y}_{x^{\prime},x}\mathcal{P}(x,y)]\psi^{y}_{x,x^{\prime}}+\sum_{x,y}\sum_{y^{\prime}}[w^{y,y^{\prime}}_{x}\mathcal{P}(x,y^{\prime})-w^{y^{\prime},y}_{x}\mathcal{P}(x,y)]\psi^{y,y^{\prime}}_{x}$
$=\sum_{x,y}\sum_{x^{\prime}}[w^{y}_{x,x^{\prime}}\mathcal{P}(x^{\prime})\mathcal{P}(y|x^{\prime})-w^{y}_{x^{\prime},x}\mathcal{P}(x)\mathcal{P}(y|x)]\psi^{y}_{x,x^{\prime}}+\sum_{x,y}\sum_{y^{\prime}}[w^{y,y^{\prime}}_{x}\mathcal{P}(x)\mathcal{P}(y^{\prime}|x)-w^{y^{\prime},y}_{x}\mathcal{P}(x)\mathcal{P}(y|x)]\psi^{y,y^{\prime}}_{x}$
$=\sum_{x,x^{\prime}}[\mathcal{P}(x^{\prime})\langle w^{y}_{x,x^{\prime}}\psi^{y}_{x,x^{\prime}}\rangle_{y|x^{\prime}}-\mathcal{P}(x)\langle w^{y}_{x^{\prime},x}\psi^{y}_{x,x^{\prime}}\rangle_{y|x}]+\Big\langle\sum_{y,y^{\prime}}[w^{y,y^{\prime}}_{x}\mathcal{P}(y^{\prime}|x)-w^{y^{\prime},y}_{x}\mathcal{P}(y|x)]\psi^{y,y^{\prime}}_{x}\Big\rangle_{x}.$
This decomposition again separates marginal and conditional contributions, in direct analogy with the FP case. For a detailed discussion of conditional corrections to EPRs see Refs. [48, 1]. As this decomposition holds for arbitrary $\bm{\psi}$, it also applies to $\langle\Delta\Sigma\rangle$ as a special case; see Eqs. (52), (78), and (97). Entropy pumping terms $\Delta\mathcal{I}$ cancel between $\Delta\mathfrak{s}_{m}$ and $\Delta\Sigma$ and thus do not contribute to the analysis of $\langle\Delta\Sigma-\Delta{\mathfrak{s}}_{m}\rangle$.
Thus, altogether, $\langle\Delta s_{\rm tot}\rangle$ satisfies properties (23)(2-3). By enforcing the second law (non-negativity) of $\langle\Delta s_{\rm tot}\rangle$, the functional is brought into full alignment with the properties (13)(1-3). Given that the KL-divergence is the unique functional satisfying the criteria of invariance, strong additivity, and non-negativity (again under mild regularity), it follows that
$\langle\Delta s_{\rm tot}\rangle\propto\lambda D_{KL},\qquad\lambda\geq 0.$ (126)
The first contribution in Eq. (118) satisfies this requirement by construction. By contrast, the second contribution contains a log ratio of conditional path weights $(\langle\Delta\mathfrak{s}_{m}\rangle)$ together with an additional term that cannot be derived from path probabilities $(\langle\Delta\Sigma\rangle)$ [82]; therefore, it does not define a KL-divergence. Consequently, imposing the second law requires $\langle\Delta\Sigma-\Delta{\mathfrak{s}}_{m}\rangle=0\Rightarrow\Delta\Sigma=\Delta{\mathfrak{s}}_{m}$. Since $\bm{\mathfrak{j}}\equiv 0$ cannot be enforced, this requirement enforces $\bm{\phi}\equiv 0$, thereby recovering LDB as a structural requirement of the second law axiom.
Appendix C Mechanical isolated system derivations
@@
In this section the detailed derivations of Sec. III.4 are provided. First, we derive the EoMs (LABEL:eq:EoM_mech_closed). The variational principle in this case is of the form:
@@
Adding both contributions then yields the total EPR given in Eq. (53).
Appendix D LDB and Master FT for closed systems
$+\frac{1}{2}(\partial_{p}F^{p}+\partial_{s}F^{s})\bigg),$ (136a)
$+\frac{1}{2}(-\partial_{p}\hat{F}^{p}+\partial_{s}\hat{F}^{s})\bigg).$ (136b)
The log ratio (11b) then gives (recall the measure contribution cancels):
which after some algebra simplifies to
Differentiating and enforcing the FDR (54), we find
$=\dot{\Sigma}.$ (139)
The last equality holds by definition, as $\dot{s}=\dot{\Sigma}$ for closed systems; see Eqs. (LABEL:eq:EoM_mech_closed). Consequently, $\Delta\Sigma$ is consistent with the log ratio of path probabilities under the constraint of the FDR, and therefore satisfies both LDB [Eq. (45)] and the master FT (III.6), as established by the argument presented in Sec. III.6. The same result holds upon inclusion of an external manipulation $\rm f$$_{+}$, demonstrating that the master FT remains valid for both closed and open systems, including those influenced by external irreversible processes or drivings.
Appendix E Interconnected systems derivations
In this section the detailed derivations of Sec. III.5 are provided. Firstly, taking variations of the Lagrangian and taking into account the fixed endpoints condition gives
@@ Appendix E Interconnected systems derivations
$\delta W^{i}:\dot{N}_{i}=(\mathcal{J}_{ji}+\sigma_{ji}\circ\zeta),$ (141d)
@@
$=\int\frac{d\Omega}{(1-C^{2})}\left\{\frac{j^{2}_{N_{1}}}{\mathcal{P}\sigma^{2}}+\frac{j^{2}_{th}}{\mathcal{P}\kappa^{2}}+2C\frac{j_{N_{1}}j_{th}}{\mathcal{P}\sigma\kappa}+j_{N_{1}}\left(\frac{\tilde{\mathcal{J}}_{12}}{\sigma^{2}}-\frac{C}{\kappa\sigma}\tilde{J}_{12}\right)-j_{th}\left(\frac{\tilde{J}_{12}}{\kappa^{2}}-\frac{C}{\kappa\sigma}\tilde{\mathcal{J}}_{12}\right)\right\},$ (145)
where we used Eqs. (74LABEL:eq:j_th_inter). Note in Sec. III.5 $\langle\!\langle\dot{\Sigma}\rangle\!\rangle,\dot{\mathcal{S}}$ are rewritten in terms of $j_{N_{2}}=-j_{N_{1}}$ for sign convenience.
For completeness, the identification of $\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle$ in Eq. (77) is verified via the PI approach. Although a brute force computation applying the Moore–Penrose pseudo-inverse to the system (71) gives the desired result, as demonstrated in Appendix D, here we follow the approach in Ref. [12], where the system (71) is rewritten as:
$\dot{N}_{1}=-\mathfrak{j}_{N},$ (146)
$\dot{N}_{2}=\mathfrak{j}_{N},$
$T^{1}\dot{s}_{1}=\mathfrak{j}_{S}(T^{1}-T^{2})+\mathfrak{j}_{N}\mu^{1},$
$T^{2}\dot{s}_{2}=\mathfrak{j}_{S}(T^{2}-T^{1})-\mathfrak{j}_{N}\mu^{2},$
$\mathfrak{j}_{N}=\mathcal{J}_{12}+\sigma_{12}\circ\zeta,$
$\mathfrak{j}_{S}=J_{12}+\kappa_{12}\circ\eta.$
Then, the PI action can be reduced to the equations for $\mathfrak{j}_{N},\mathfrak{j}_{S}$ which give the non-degenerate covariance matrix $\mathbb{L}$. Note this structure is naturally revealed by the Shannon entropy change in Eq. (77). The forward and backward actions then take the form:
$\mathcal{A}[\{N_{i}\},\{s_{i}\}]=\int_{t_{1}}^{t_{2}}\frac{d\tau}{4}\begin{bmatrix}\mathfrak{j}_{N}-\tilde{\mathcal{J}}_{12}\\ \mathfrak{j}_{S}-\tilde{J}_{12}\end{bmatrix}\mathbb{L}^{-1}\begin{bmatrix}\mathfrak{j}_{N}-\tilde{\mathcal{J}}_{12}\\ \mathfrak{j}_{S}-\tilde{J}_{12}\end{bmatrix}^{T},$ (147)
$\hat{\mathcal{A}}[\{N_{i}\},\{s_{i}\}]=\int_{t_{1}}^{t_{2}}\frac{d\tau}{4}\begin{bmatrix}-\mathfrak{j}_{N}-\tilde{\mathcal{J}}_{12}\\ -\mathfrak{j}_{S}-\tilde{J}_{12}\end{bmatrix}\mathbb{L}^{-1}\begin{bmatrix}-\mathfrak{j}_{N}-\tilde{\mathcal{J}}_{12}\\ -\mathfrak{j}_{S}-\tilde{J}_{12}\end{bmatrix}^{T},$ (148)
where $\mathfrak{j}_{N},\mathfrak{j}_{S}$ are odd under time-reversal (see [12]). After some algebra it is found:
$\Delta\mathfrak{s}_{m}=\hat{\mathcal{A}}-\mathcal{A}=\int_{t_{1}}^{t_{2}}d\tau\begin{bmatrix}\tilde{\mathcal{J}}_{12}\\ \tilde{J}_{12}\end{bmatrix}\mathbb{L}^{-1}\begin{bmatrix}\mathcal{J}_{12}+\sigma_{12}\circ\zeta\\ J_{12}+\kappa_{12}\circ\eta\end{bmatrix}^{T}.$ (149)
Taking the time derivative followed by the two-step average yields:
$\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle=\int d\Omega\begin{bmatrix}j_{N_{2}}\\ j_{th}\end{bmatrix}^{T}\mathbb{L}^{-1}\begin{bmatrix}\tilde{\mathcal{J}}_{12}\\ \tilde{J}_{12}\end{bmatrix}$, (150)
in agreement with Eq. (77).
Appendix F Mechanical open system derivations
In this section the detailed derivations of Sec. III.6 are provided. First, we derive the EoMs (LABEL:eq:EoM_open). The variational principle in this case is of the form: